Required information [The following information applies to the questions displayed below.) Bunnell Corporation is a manufactu12. What is the ending balance in Finished Goods? Finished Goods Beg. Bal. End. Bal.13. Assuming that the company closes its underapplied or overapplied overhead to Cost of Goods Sold, what is the adjusted cos14. What is the gross margin for the year? Gross margin15. What is the net operating income for the year? Net operating income
Answer #1

12) Finished goods

Beg bal 36000 Cost of goods sold 1488700
Cost of goods manufactured 1478700
End bal 26000

13) Applied overhead = 41000*13 = 533000

Actual overhead = 150000+370000 = 520000

Over applied overhead = 13000

Adjusted cost of goods sold = 1488700-13000 = 1475700

14) Gross margin = 3367500-1475700 = 1891800

15) Net operating income = 1891800-249000-464000 = 1178800